# Trees of Graphs as Boundaries of Hyperbolic Groups

### v1

@@ 9.2. The extended Whitehead graph $\widetilde{W}_{v}$ of a virtually free rigid cluster factor $G(v)$
It remains to describe the set $\mathcal{A}$ of $E$-connections for $\Gamma$ appearing in ${\mathcal{R}}_{G}$. For this we need some preparations which concern the action of $G$ on the tree $T^{r}_{G}$. For each vertex $z\in Z$ of type (v2), fix any bijective correspondence between the set of edges of the $\theta$-graph $\Gamma_{z}=\Xi_{k(z)}$ and the set $N_{z}$ of vertices of $T^{r}_{G}$ adjacent to $z$ (which are all flexible). Denote this bijection by $\lambda_{z}^{+}$. Fix also an identification of the vertex set of $\Gamma_{z}$ (which consists of 2 vertices) with the doubleton $\partial G(z)\subset\partial G$. Next, recall that for each rigid cluster vertex $z\in Z$, we have a natural bijective correspondence between the set of orbits in $N_{z}$ under the action of $G(z)$ (i.e. the quotient $N_{z}/G(z)$) and the set ${\mathcal{H}}_{z}$ (of conjugacy classes of peripheral subgroups of $G(z)$, under the view of $G(z)$ as an abstract factor, see Example 8.2). This correspondence is given by assigning to any orbit $[u]\in N_{z}/G(z)$ (with $u\in N_{z}$ as any representative of this orbit) the conjugacy class of the subgroup $G(u)<G(z)$ (note that, due to Lemma 9.1, the groups $G(u)$ are indeed contained in $G(z)$). For any $\xi\in{\mathcal{H}}_{z}$, we fix some representative $u_{\xi}\in N_{z}$ of the orbit in $N_{z}/G(z)$ corresponding to $\xi$, and we also fix a bijective correspondence between the vertices in the set $N_{u_{\xi}}\setminus\{z\}$ (which are all flexible, by the fact that $z$ is isolated, see Fact 3.5) and the set of numbers $11,2,\dots,n_{\xi}-$ 1, where $n_{\xi}$ is the degree of $u_{\xi}$ in $T^{r}_{G}$. This induces a bijective correspondence between the set ~$\widetilde{\mathcal{H}}_{z}=\{(\xi,j):\xi\in{\mathcal{H}}_{z},1\leq j\leq n_{\xi}-1\}$ and the set $N^{2}_{z}$ of all those vertices $y$ in $T^{r}_{G}$ lying at distance 2 from $z$, for which the shortest path between $y$ and $z$ in $T^{r}_{G}$ passes through one of the vertices $u_{\xi}$. We denote by $\lambda^{+}_{z}:E_{\widetilde{W}_{z}}\to N^{2}_{z}$ the labelling $\tilde{\lambda}_{z}$ in which the labels from ~$\widetilde{\mathcal{H}}_{z}$ are replaced with the corresponding elements from $N^{2}_{z}$ (where $\tilde{\lambda}_{z}$ is the labelling described in Section 9, right before Fact 9.5).
@@
To describe the set $\mathcal{A}$, consider the set $Y$ of orbits of the action of $G$ on the set of all flexible vertices of $T^{r}_{G}$. The set $\mathcal{A}$ will be given as the (disjoint) union of the sets ${\mathcal{A}}_{y}:y\in Y$, which we call blocks. We now pass to describing these blocks ${\mathcal{A}}_{y}$. Fix any orbit $y\in Y$. For any $z\in Z$ of type $(v2)$, let $y^{z}_{i}:i\in I_{y,z}$ be the set of all vertices adjacent to $z$ in the orbit $y$. Similarly, for any rigid cluster vertex $z\in Z$, let $y^{z}_{i}:i\in I_{y,z}$ be the set of all vertices from the set $N^{2}_{z}$ in the orbit $y$. (Note that for some of $z$ the index sets $I_{y,z}$ may be empty.) Now, if $z\in Z$ is of type (v2), each $y_{i}^{z}$ as above belongs to $N_{z}$, and hence the preimage $(\lambda^{+}_{z})^{-1}(y_{i}^{z})$ is a subset of the edge set of $\Gamma_{z}=\Xi_{k(z)}$ consistsing of a single edge, namely the edge labelled with $y_{i}^{z}$. If $z\in Z$ is a rigid cluster vertex, each $y_{i}^{z}$ as above belongs to the set $N^{2}_{z}$, and hence the preimage $(\lambda^{+}_{z})^{-1}(y_{i}^{z})$ is a subset of the edge set of the graph $\Gamma_{z}=\widetilde{W}_{z}$, namely it also consists of these edges of $\widetilde{W}_{z}$ which are labelled with $y_{i}^{z}$ (this time there may be more than one such edge). Put
@@
It is not hard to observe that the family $E_{y}:y\in Y$ is then a partition of the edge set of $\Gamma$, i.e. each edge of $\Gamma$ belongs to precisely one of the sets $E_{y}$.
@@
Case 1. If $z$ is a rigid cluster vertex (so that $\Gamma_{z}$ is the extended Whitehead graph for the rigid cluster factor ${\mathcal{Q}}_{z}$), put $\xi=\lambda_{z}(\varepsilon)\in{\mathcal{H}}_{z}$. We then have $y_{\varepsilon}\in N_{u_{\xi}}\setminus\{z\}$, so the fixed orientation of $G(y_{\varepsilon})$ induces an orientation of its peripheral subgroup $G(y_{\varepsilon})\cap G(u_{\xi})$. This obviously determines an order of the doubleton $\partial G(u_{\xi})$. Since $G(u_{\xi})$ is a peripheral subgroup of the factor $G(z)$ (compare Lemma 9.1), $\partial G(u_{\xi})$ is a peripheral of the associated space with peripherals for $G(z)$. Identifying the latter naturally with the core $C(\widetilde{\mathcal{V}}_{z})$ of the associated $V$-tree $X_{V}(\widetilde{\mathcal{V}}_{z})$ (as in Lemma 9.6(1)), $\partial G(u_{\xi})$ is a peripheral of this core. Since $\varepsilon$ is an edge of the extended Whitehead graph $\widetilde{W}_{z}=\Gamma_{z}$, and since it belongs to edges associated to the peripheral $\partial G(u_{\xi})$, it carries the order-induced orientation (with respect to the above described order of $\partial G(u_{\xi})$). We denote by $\varepsilon^{\#}$ the edge $\varepsilon$ with this order-induced orientation.
@@
Having defined all the blocks ${\mathcal{A}}_{y}$, we put ${\mathcal{A}}=\bigcup_{y\in Y}{\mathcal{A}}_{y}$. We skip the straightforward verification of the fact that the so described set ${\mathcal{A}}$ does not depend on the choices of the set of representatives $Z$, and the representatives $v_{y}$ of the orbits $y\in Y$. This completes the description of the graphical connecting system ${\mathcal{R}}_{G}$. The fact that it satisfies the transitivity property (4) from Definition 6.1 follows, roughly speaking, from the fact that the tree $T^{r}_{G}$ is connected, and we omit the details.
Given the graphical connecting system ${\mathcal{R}}_{G}$, as introduced above, we may state the following result, which describes explicitely the form of a regular tree of graphs which appears only imlicitely in our statement of Therorem A (in conditions (2) and (3)). This result, Theorem 10.1 below, may be viewed as the third main result of the paper. Its proof, after necessary preparations provided in the next two sections, will be presented in Section 13.
@@
Let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian, and suppose that each rigid cluster factor of its reduced Bowditsch’s JSJ splitting is virtually free. Let ${\mathcal{R}}_{G}$ be the graphical connecting system associated to $G$ (as described above in this section). Then the Gromov boundary $\partial G$ is homeomorphic to the regular tree of graphs ${\mathcal{X}}({\mathcal{R}}_{G})$.
@@
Now, by the above description of ${\mathcal{R}}_{G}$, any bunch of edges $\mathcal{M}$ labelled with $\xi$ is in a fixed bijective correspondence (provided by the labelling $\lambda_{z}^{+}$) with the vertex set $N_{u_{\xi}}\setminus\{z\}$. This correspondence allows to split $\mathcal{M}$ into subsets corresponding to the orbits of the action of $G(u_{\xi})$ on $N_{u_{\xi}}\setminus\{z\}$. More precisely, observe that due to Lemma 9.1, $z$ is a global fixed point of the action of $G(u_{\xi})$ on $N_{u_{\xi}}$ (and an alternative justification follows by the fact that $z$ is the unique rigid cluster vertex in $N_{u_{\xi}}$, due to its isolatedness). Consequently, the $G(u_{\xi})$-orbit $[x]$ of any vertex $x\in N_{u_{\xi}}\setminus\{z\}$ is contained in $N_{u_{\xi}}\setminus\{z\}$. The subset of $\mathcal{M}$ corresponding to any such orbit $[x]$ will be called a packet of edges in $\widetilde{W}_{z}$.
We will later need the following easy but rather technical observation, the proof of which we omit.
@@
    Having fixed $\mathcal{M}$, for all $M$ and $x$ as above for which the corresponding flexible factors $G(x)$ are orientable, the orientations of all (oriented) edges $\varepsilon^{\#}$ with $\varepsilon$ in any such $M$ (as defined in the above description of the set of E-connections for ${\mathcal{R}}_{G}$) are compatible, i.e. they induce the same order on the common set of endpoints of these edges.
  3. (3)
    For any fixed $M$ and $x$ as in the assumption, if $G(x)$ is non-orientable (as a flexible factor), and if the edge group $G([u_{\xi},x])$ is orientable, then the cardinality of the packet $M$ is even, and half of the orientations of the edges $\{\varepsilon^{\#}:\varepsilon\in M\}$ are compatible with one of the orders in the common set of endpoints, while the other half of these orientations are compatible with the opposite order.
@@
    For all $M$ and $x$ as above for which the corresponding flexible factors $G(x)$ are orientable, the orientations of all (oriented) edges $\varepsilon^{\#}$ with $\varepsilon$ in any such $M$ (as defined in the above description of the set of E-connections for ${\mathcal{R}}_{G}$) are compatible, i.e. they induce the same order on the common set of endpoints of these edges.
  3. (3)
    For any fixed $M$ and $x$ as in the assumption, if $G(x)$ is non-orientable (as a flexible factor), and if the edge group $G([u_{\xi},x])$ is orientable, then the cardinality of the packet $M$ is even, and half of the orientations of the edges $\{\varepsilon^{\#}:\varepsilon\in M\}$ are compatible with one of the orders in the common set of endpoints, while the other half of these orientations are compatible with the opposite order.
@@ 11. Regular tree systems of spaces with doubleton peripherals
In this section we prepare some tools and terminology needed in our proof of Theorem 10.1 (which gives also implications (3$\Rightarrow$(2) in Theorem A and (3$\Rightarrow$(2) in Theorem B). Recall that a space with doubleton peripherals is a space with peripherals in which each peripheral subspace consists of precisely two points (i.e. is a doubleton). We start with a rather long definition.
@@
We now present few classes of examples of regular spaces with doubleton peripherals. These examples will be used in our later considerations and arguments concerning the proof of Theorem 10.1.
@@
Recall that to any abstract factor ${\mathcal{F}}=(K,\{H_{i}\}_{i\in I})$ (as in Section 8.1) there is associated the induced space with doubleton peripherals, as described in Definition 8.3; we denote this space with peripherals by $X_{\mathcal{F}}$. We describe a decoration ${\mathcal{D}}={\mathcal{D}}_{\mathcal{F}}$ naturally associated to $X_{\mathcal{F}}$, which turns out to satisfy the regularity condition of Definition 11.1(4), and which yields what we call the regular spaces with doubleton peripherals associated to an abstract factor.
The labelling $\tau$ of $\mathcal{D}$ is given by any bijection $b:I\to{\mathcal{T}}$ through the formula
$\tau(\partial(gH_{i}g^{-1})):=b(i).$
The orientability (and non-orientability) data for the peripherals, as well as the assigned orientations, are described by inducing in the obvious way the similar data for the corresponding peripheral subgroups (provided by a fixed orientation data for the factor, as given in Definition 8.5). By Lemma 8.6, the so described decoration ${\mathcal{D}}_{\mathcal{F}}$ satisfies the regularity condition.
@@
Let $(K,\Omega)$ be a punctured circle, i.e. a space with peripherals (as described right before Definition 2.3) in which $K=(S^{1})^{\circ}$, with its standard family $\Omega$ of peripheral subspaces, which are doubletons. Fix an arbitrary auxiliary orientation of $(K,\Omega)$ (in the sense of Remark 8.9). Let ${\mathcal{D}}$ be any decoration for $(K,\Omega)$ which is dense, i.e. satisfies one of the following two conditions:
@@
    for each type of peripherals, the union of all (underlying spaces of) directed peripherals of this type in $((K,\Omega),{\mathcal{D}})$, with orientations compatible (respectively, incompatible) with the fixed orientation of $(K,\Omega)$, is dense in $K$; less formally, if a type of peripherals is non-orientable, we simply require that the union of peripherals of this type is dense in $K$, and if a type in orientable, we require that both unions of all those peripherals of this type whose orientations are compatible (respectively, incompatible) with the fixed orientation of $(K,\Omega)$ are dense in $K$;
@@
We then say that the pair $((K,\Omega),{\mathcal{D}})$ is a densely decorated punctured circle. We will speak of two kinds of densely decorated punctured circles: those satisfying condition (a) above will be called non-orinetable, and those satisying condition (b) - orientable. Next lemma collects some usefull properties of densely decorated punctured circles, and we omit its straightforward proof which can be performed by referring to the standard back and forth argument.
@@
    Any two orientable densely decorated punctured circles with the same number of types of peripherals are isomorhic (via a decoration-preserving homeomorhism $h$), after appropriate identification of the types, and after possibly reversing orientations of all peripherals of some of the types. Moreover, for any bijection $\beta$ between the sets of types of the punctured circles above (which we view as the identification of the types) there is a homeomorphism $h$ as above which is compatible with $\beta$ (i.e. such that for any peripheral $P$ we have $\tau(h(P))=\beta(\tau(P))$).
@@
    Any two non-orientable densely decorated punctured circles with the same number of types of orientable peripherals, and the same number of types of non-orientable peripherals, are isomorhic, after appropriate identification of the types. Moreover, for any bijection $\beta$ between the sets of types of the punctured circles above, if $\beta$ respects orientability and non-orientability of the types, then there is an isomorphism $h$ between those decorated punctured circles which is compatible with $\beta$.
@@
We now proceed one level further, and discuss some tree systems made out of regular spaces with doubleton peripherals. As we will see, such tree systems are induced, in an essentially unique way, by a rather simple data described in the next definitions.
@@
Given an arbitrary regular (decorated) space with doubleton peripherals $((K,\Omega),{\mathcal{D}})$, a degree function for this space is a positive integer valued function $d:\Omega\to Z$ which is constant on types, i.e. such that $d(P_{1})=d(P_{2})$ for any two peripherals $P_{1},P_{2}$ of the same type. We call a collection of data ${\mathcal{D}}^{+}=({\mathcal{D}},d)$ consisting of a decoration $\mathcal{D}$ and a degree function $d$ an extra-decoration for $(K,\Omega)$. The pair $((K,\Omega),{\mathcal{D}}^{+})$ is then called an extra-decorated regular space with doubleton peripherals.
@@
    for each block $B_{a}$, an additional information whether $B_{a}$ is orientable or non-orientable; we require that if $B_{a}$ is orientable then it necessarily consists of orientable types only, but if $B_{a}$ is non-orientable, there is no restriction on orientability and non-orientability of the types that it contains.
We also require in the definition that the block structure $\mathcal{B}$ fulfills the following transitivity condition: for any two indices $j,j^{\prime}\in J$ there is a sequence of indices $j=j_{0},j_{1},\dots,j_{k}=j^{\prime}$ such that for each $i=1,\dots,k$ there exist peripherals $P\in\Omega_{j_{i-1}}$ and $P^{\prime}\in\Omega_{j_{i}}$ whose types belong to the same block.

### v2

@@ 9.2. The extended Whitehead graph $\widetilde{W}_{v}$ of a virtually free rigid cluster factor $G(v)$
It remains to describe the set $\mathcal{A}$ of $E$-connections for $\Gamma$ appearing in ${\mathcal{R}}_{G}$. For this we need some preparations which concern the action of $G$ on the tree $T^{r}_{G}$. For each vertex $z\in Z$ of type (v2), fix any bijective correspondence between the set of edges of the $\theta$-graph $\Gamma_{z}=\Xi_{k(z)}$ and the set $N_{z}$ of vertices of $T^{r}_{G}$ adjacent to $z$ (which are all flexible). Denote this bijection by $\lambda_{z}^{+}$. Fix also an identification of the vertex set of $\Gamma_{z}$ (which consists of 2 vertices) with the doubleton $\partial G(z)\subset\partial G$. Next, recall that for each rigid cluster vertex $z\in Z$, we have a natural bijective correspondence between the set of orbits in $N_{z}$ under the action of $G(z)$ (i.e. the quotient $N_{z}/G(z)$) and the set ${\mathcal{H}}_{z}$ (of conjugacy classes of peripheral subgroups of $G(z)$, under the view of $G(z)$ as an abstract factor, see Example 8.2). This correspondence is given by assigning to any orbit $[u]\in N_{z}/G(z)$ (with $u\in N_{z}$ as any representative of this orbit) the conjugacy class of the subgroup $G(u)<G(z)$ (note that, due to Lemma 9.1, the groups $G(u)$ are indeed contained in $G(z)$). For any $\xi\in{\mathcal{H}}_{z}$, we fix some representative $u_{\xi}\in N_{z}$ of the orbit in $N_{z}/G(z)$ corresponding to $\xi$, and we also fix a bijective correspondence between the vertices in the set $N_{u_{\xi}}\setminus\{z\}$ (which are all flexible, by the fact that $z$ is isolated, see Fact 3.5) and the set of numbers $11,2,\dots,n_{\xi}-$ 1, where $n_{\xi}$ is the degree of $u_{\xi}$ in $T^{r}_{G}$. This induces a bijective correspondence between the set ~$\widetilde{\mathcal{H}}_{z}=\{(\xi,j):\xi\in{\mathcal{H}}_{z},1\leq j\leq n_{\xi}-1\}$ and the set $N^{2}_{z}$ of all those vertices $y$ in $T^{r}_{G}$ lying at distance 2 from $z$, for which the shortest path between $y$ and $z$ in $T^{r}_{G}$ passes through one of the vertices $u_{\xi}$. We denote by $\lambda^{+}_{z}:E_{\widetilde{W}_{z}}\to N^{2}_{z}$ the labelling obtained from $\tilde{\lambda}_{z}$ (where $\tilde{\lambda}_{z}$ is the labelling described in Section 9, right before Fact 9.5) by replacing the labels from ~$\widetilde{\mathcal{H}}_{z}$ with the corresponding elements from $N^{2}_{z}$.
@@
To describe the set $\mathcal{A}$, consider the set $Y$ of orbits of the action of $G$ on the set of all flexible vertices of $T^{r}_{G}$. The set $\mathcal{A}$ will be given as the (disjoint) union of the sets ${\mathcal{A}}_{y}:y\in Y$, which we call blocks. We now pass to describing these blocks ${\mathcal{A}}_{y}$. Fix any orbit $y\in Y$. For any $z\in Z$ of type $(v2)$, let $y^{z}_{i}:i\in I_{y,z}$ be the set of all vertices adjacent to $z$ in the orbit $y$. Similarly, for any rigid cluster vertex $z\in Z$, let $y^{z}_{i}:i\in I_{y,z}$ be the set of all vertices from the set $N^{2}_{z}$ in the orbit $y$. (Note that for some $z$, the index sets $I_{y,z}$ may be empty.) Now, if $z\in Z$ is of type (v2), each $y_{i}^{z}$ as above belongs to $N_{z}$, and hence the preimage $(\lambda^{+}_{z})^{-1}(y_{i}^{z})$ is a subset of the edge set of $\Gamma_{z}=\Xi_{k(z)}$ consistsing of a single edge, namely the edge labelled with $y_{i}^{z}$. If $z\in Z$ is a rigid cluster vertex, each $y_{i}^{z}$ as above belongs to the set $N^{2}_{z}$, and hence the preimage $(\lambda^{+}_{z})^{-1}(y_{i}^{z})$ is a subset of the edge set of the graph $\Gamma_{z}=\widetilde{W}_{z}$, namely it also consists of these edges of $\widetilde{W}_{z}$ which are labelled with $y_{i}^{z}$ (this time there may be more than one such edge). Put
@@
It is not hard to observe that the family $E_{y}:y\in Y$ is then a partition of the edge set of $\Gamma$, i.e., each edge of $\Gamma$ belongs to precisely one of the sets $E_{y}$.
@@
Case 1. If $z$ is a rigid cluster vertex (so that $\Gamma_{z}$ is the extended Whitehead graph for the rigid cluster factor ${\mathcal{Q}}_{z}=(G(z),\mathcal{H}_{z})$), put $\xi=\lambda_{z}(\varepsilon)\in{\mathcal{H}}_{z}$. We then have $y_{\varepsilon}\in N_{u_{\xi}}\setminus\{z\}$, so the fixed orientation of $G(y_{\varepsilon})$ induces an orientation of its peripheral subgroup $G(y_{\varepsilon})\cap G(u_{\xi})$. This obviously determines an order of the doubleton $\partial G(u_{\xi})$. Since $G(u_{\xi})$ is a peripheral subgroup of the factor $G(z)$ (compare Lemma 9.1), $\partial G(u_{\xi})$ is a peripheral of the space with peripherals associated to $\mathcal{Q}_{z}$. Identifying the latter naturally with the core $C(\widetilde{\mathcal{V}}_{z})$ of the associated $V$-tree ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{v})$ (as in Lemma 9.6(1)), $\partial G(u_{\xi})$ is a peripheral of this core. Since $\varepsilon$ is an edge of the extended Whitehead graph $\widetilde{W}_{z}=\Gamma_{z}$, and since it belongs to edges associated to the peripheral $\partial G(u_{\xi})$, it carries the order-induced orientation (with respect to the above described order of $\partial G(u_{\xi})$). (See Remark 9.10(3).) We denote by $\varepsilon^{\#}$ the edge $\varepsilon$ with this order-induced orientation.
@@
Having defined all the blocks ${\mathcal{A}}_{y}$, we put ${\mathcal{A}}=\bigcup_{y\in Y}{\mathcal{A}}_{y}$. We skip the straightforward verification of the fact that the so described set ${\mathcal{A}}$ does not depend on the choices of the set of representatives $Z$, and the representatives $v_{y}$ of the orbits $y\in Y$. This completes the description of the graphical connecting system ${\mathcal{R}}_{G}$. The fact that ${\mathcal{R}}_{G}$ satisfies the transitivity property (4) from Definition 6.1 follows, roughly speaking, from the fact that the tree $T^{r}_{G}$ is connected, and we omit the details.
Given the graphical connecting system ${\mathcal{R}}_{G}$, as introduced above, we may state the following result, which describes explicitly the form of a regular tree of graphs which appears only implicitly in our statement of Therorem A (in condition (2)). This result, Theorem 10.1 below, may be viewed as the third main result of the paper. Its proof, after necessary preparations provided in the next two sections, will be presented in Section 13.
@@
Let $G$ be a 1-ended hyperbolic group which is not cocompact Fuchsian, and suppose that each rigid cluster factor of its reduced Bowditch JSJ splitting is virtually free. Let ${\mathcal{R}}_{G}$ be the graphical connecting system associated to $G$ (as described above in this section). Then the Gromov boundary $\partial G$ is homeomorphic to the regular tree of graphs ${\mathcal{X}}({\mathcal{R}}_{G})$.
@@
Now, as mentioned in the above description of ${\mathcal{R}}_{G}$, any bunch of edges $\mathcal{M}$ labelled with $\xi$ is in a fixed bijective correspondence (provided by the labelling $\lambda_{z}^{+}$) with the vertex set $N_{u_{\xi}}\setminus\{z\}$. This correspondence allows us to split $\mathcal{M}$ into subsets corresponding to the orbits of the action of $G(u_{\xi})$ on $N_{u_{\xi}}\setminus\{z\}$. More precisely, observe that due to Lemma 9.1, $z$ is a global fixed point of the action of $G(u_{\xi})$ on $N_{u_{\xi}}$ (and an alternative justification follows by the fact that $z$ is the unique rigid cluster vertex in $N_{u_{\xi}}$, due to its isolatedness). Consequently, the $G(u_{\xi})$-orbit $[x]$ of any vertex $x\in N_{u_{\xi}}\setminus\{z\}$ is contained in $N_{u_{\xi}}\setminus\{z\}$. The subset of $\mathcal{M}$ corresponding to any such orbit $[x]$ will be called a packet of edges in $\widetilde{W}_{z}$.
We will later need the following easy but a bit technical observation, the proof of which we omit.
@@
    Having fixed $\mathcal{M}$, for all $M$ and $x$ as above for which the corresponding flexible factor $G(x)$ is orientable, we have
    1. (a)
      if the group $G(u_{\xi})$ is orientable then the orientations of all (oriented) edges $\varepsilon^{\#}$ with $\varepsilon$ in $M$ (as defined in the above description of the set of E-connections for ${\mathcal{R}}_{G}$) are compatible, i.e., they induce the same order on the common set of endpoints of these edges;
    2. (b)
      if the group $G(u_{\xi})$ is non-orientable then the cardinality of the packet $M$ is even, and half of the orientations of the edges $\{\varepsilon^{\#}:\varepsilon\in M\}$ are compatible with one of the orders in the common set of endpoints, while the other half of these orientations are compatible with the opposite order.
@@
    For all $M$ and $x$ as above for which the corresponding flexible factor $G(x)$ is orientable, we have
    1. (a)
      if $G(z)$ is orientable then the orientations of all (oriented) edges $\varepsilon^{\#}$ with $\varepsilon$ in any such $M$ (as defined in the above description of the set of E-connections for ${\mathcal{R}}_{G}$) are compatible, i.e., they induce the same order on the common set of endpoints of these edges;
    2. (b)
      if $G(z)$ is non-orientable then the cardinality of the packet $M$ is even, and half of the orientations of the edges $\{\varepsilon^{\#}:\varepsilon\in M\}$ are compatible with one of the orders in the common set of endpoints, while the other half of these orientations are compatible with the opposite order.
@@ 11. Regular tree systems of spaces with doubleton peripherals
In this section we prepare some tools and terminology needed in our proof of Theorem 10.1 (from which we then deduce the implications (3$\Rightarrow$(2) in Theorem A and (3$\Rightarrow$(2) in Theorem B). By a space with doubleton peripherals we mean a space with peripherals in which each peripheral subspace consists of precisely two points (i.e. is a doubleton). We start with a rather long definition.
@@
We now present a few classes of examples of regular spaces with doubleton peripherals. These examples will be used in our later considerations and arguments concerning the proof of Theorem 10.1.
@@
Recall that to any abstract factor ${\mathcal{F}}=(K,\{H_{i}\}_{i\in I})$ (as in Section 8.1) there is associated the induced space with doubleton peripherals, as described in Definition 8.3; we denote this space with peripherals by $X_{\mathcal{F}}$. We describe a decoration ${\mathcal{D}}={\mathcal{D}}_{\mathcal{F}}$ naturally associated to $X_{\mathcal{F}}$, which turns out to satisfy the regularity condition of Definition 11.1(4), and which yields what we call the regular space with doubleton peripherals associated to an abstract factor.
The labelling $\tau$ of $\mathcal{D}$ has values in the set ${\mathcal{T}}:=I$ and is given by the formula
$\tau(\partial(gH_{i}g^{-1})):=i.$
The orientability (and non-orientability) data for the peripherals, as well as the assigned orientations, are described by inducing in the obvious way the similar data for the corresponding peripheral subgroups (provided by fixed orientation data for the factor, as given in Definition 8.5). By Lemma 8.6, the so described decoration ${\mathcal{D}}_{\mathcal{F}}$ satisfies the regularity condition.
@@
Let $(K,\Omega)$ be a punctured circle, i.e., a space with peripherals (as described right before Definition 2.3) in which $K=(S^{1})^{\circ}$, with its standard family $\Omega$ of peripheral subspaces, which are doubletons. Fix an arbitrary auxiliary orientation of $(K,\Omega)$ (in the sense of Remark 8.9). Let ${\mathcal{D}}$ be any decoration for $(K,\Omega)$ which is dense, i.e., satisfies one of the following two conditions:
@@
    for each type of peripheral, the union of all directed peripherals of this type in $((K,\Omega),{\mathcal{D}})$, with orientations compatible (respectively, incompatible) with the fixed orientation of $(K,\Omega)$, is dense in $K$; that is, if a type of peripherals is non-orientable, we require simply that the union of peripherals of this type is dense in $K$, and if a type is orientable, we require both that the union of all those peripherals of this type whose orientations are compatible with the fixed orientation of $(K,\Omega)$ is dense in $K$ and that the union of all those peripherals of this type whose orientations are incompatible with the fixed orientation of $(K,\Omega)$ is dense in $K$;
@@
We then say that the pair $((K,\Omega),{\mathcal{D}})$ is a densely decorated punctured circle. We will speak of two kinds of densely decorated punctured circles: those satisfying condition (a) above will be called non-orientable, and those satisying condition (b) will be called orientable. The next lemma collects some useful properties of densely decorated punctured circles, and we omit its straightforward proof which can be performed by a standard back and forth argument.
@@
    Any two orientable densely decorated punctured circles with the same number of types of peripherals are isomorphic (via a decoration-preserving homeomorhism $h$), after appropriate identification of the types, and after possibly reversing orientations of all peripherals of some of the types. Moreover, for any bijection $\beta$ between the sets of types of the punctured circles above (which we view as the identification of the types) there is a homeomorphism $h$ as above which is compatible with $\beta$ (i.e. such that for any peripheral $P$ we have $\tau(h(P))=\beta(\tau(P))$).
@@
    Any two non-orientable densely decorated punctured circles with the same number of types of orientable peripherals, and the same number of types of non-orientable peripherals, are isomorphic, after appropriate identification of the types. Moreover, for any bijection $\beta$ between the sets of types of the punctured circles above, if $\beta$ respects orientability and non-orientability of the types, then there is an isomorphism $h$ between those decorated punctured circles which is compatible with $\beta$.
@@
We now proceed one level further, and discuss some tree systems made out of regular spaces with doubleton peripherals. As we will see, such tree systems are induced, in an essentially unique way, by rather simple data described in the next definitions.
@@
Given an arbitrary regular (decorated) space with doubleton peripherals $((K,\Omega),{\mathcal{D}})$, a degree function for this space is a positive integer valued function $d:\Omega\to\mathbb{Z}$ which is constant on types, i.e., such that $d(P_{1})=d(P_{2})$ for any two peripherals $P_{1},P_{2}$ of the same type. We call a collection of data ${\mathcal{D}}^{+}=({\mathcal{D}},d)$ consisting of a decoration $\mathcal{D}$ and a degree function $d$ an extra-decoration for $(K,\Omega)$. The pair $((K,\Omega),{\mathcal{D}}^{+})$ is then called an extra-decorated regular space with doubleton peripherals.
@@
    for each block $B_{a}$, additional on information whether $B_{a}$ is orientable or non-orientable; we require that if $B_{a}$ is orientable then it necessarily consists of orientable types only, but if $B_{a}$ is non-orientable, there is no restriction on orientability and non-orientability of the types that it contains.
We also require in the definition that the block structure $\mathcal{B}$ fulfills the following transitivity condition: for any two indices $j,j^{\prime}\in J$ there is a sequence of indices $j=j_{0},j_{1},\dots,j_{k}=j^{\prime}$ such that for each $i=1,\dots,k$ there exist types $\tau\in{\mathcal{T}}_{j_{i-1}}$ and $\tau^{\prime}\in{\mathcal{T}}_{j_{i}}$ which belong to the same block.